A truck covers 40.0 m in 9.45 s while uniformly slowing down to a final velocity of 2.35 m/s.

(a) Find the truck's original speed.
________m/s

(b) Find its acceleration.
_________ m/s2

A car starts from rest and travels for 7.5 s with a uniform acceleration of +3.0 m/s2. The driver then applies the brakes, causing a uniform acceleration of −1.5 m/s2. If the brakes are applied for 2.0 s, determine each of the following.

(c) How fast is the car going at the end of the braking period?
m/s

(d) How far has the car gone?
m
Answer #1

u a = +3.0 m/s2 a = -1.5m/sh initial velocity u= 0 u= us at = 3.0x 7.5 = 22.5 m/s is then acceleration a = – 1.5 mlr V2 = vtu a = +3.0 m/s2 a = -1.5m/sh initial velocity u= 0 u= us at = 3.0x 7.5 = 22.5 m/s is then acceleration a = – 1.5 mlr V2 = vtu²+baa = 2.352 t=9.45 solution :- &= 40 m, V=2.35 mls. final velocity 22= 42 + 2 as, 2.35² = 42 + 2.9x40 and ut at a x 9.45 2